OpenStudy (anonymous):

I'm trying to find the field lines, r=(x(t),y(t),z(t)), of F(x,y,z)=(y-z,-x,x-y). Am I supposed to use the formula: \[\dot{\underline{r}}(t)=\lambda(\underline{r}(t))\underline{F}(\underline{r}(t))\] for some scalarfield lambda? How does it work?

OpenStudy (anonymous):

F(x,y,z)=(y-z,z-x,x-y)

OpenStudy (anonymous):

I don't recognize this formula, is it some kind of graphical way to represent div and curl of a vector field?

OpenStudy (anonymous):

\[\dot{\underline{r}}(t)=(\frac{dx(t)}{dt},\frac{dy(t)}{dt},\frac{dz(t)}{dt})\] So I think you're supposed to get a system of differential equations. The line underneath letters means that they're vectors by the way. If you have a different formula, I'm open to ideas.